OpenStudy (anonymous):

can anyone explain a for statement in vb.net for me? my instructor isn't doing that great of a job. :/

OpenStudy (anonymous):

Try searching youtube for some of those tutorial videos (specifically on Visual Basic .NET).

OpenStudy (anonymous):

I don't know visual basic .net, but the general idea of a for statement is to enter a loop. pseudo-code: for (start when i = 0, while i < # of iterations you need, increase i after each iteration) { print hello world } example: for( i=0, i<5, i++) { print hello } so the iterations will look like: is 0<5? yes, i = 0, hello... i=0+1 is 1<5? yes, i = 1, hello... i=1+1 is 2<5? yes, i = 2, hello... i=2+1 is 3<5? yes, i =3, hello... i =3+1 is 4<5? yes, i = 4, hello... i=4+1 is 5<5? no. end for loop. At the end of the loop, 5 hellos are printed. I hope this helps some.